OpenStudy (anonymous):

\[f(x)= \frac{ ax ^{3}+bx^{2}-13x+2c }{ x^{2}-2x+1 }, and \lim_{x \rightarrow 1}f(x) = 11\] Find a,b,c

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@AravindG

hartnn (hartnn):

since the limit exist, x-1 is the factor of numerator so 1 is the root.--->a+b+2c-13=0 since u have a factor of x-1 twice in the denominator, u need to use L'Hopitals twice and put x=1. so u have 3a+2b-13=0 and lastly, 6a+b=11 solve these three simultaneously, u'll find a,b,c.

OpenStudy (anonymous):

right thanks

hartnn (hartnn):

welcome :)

OpenStudy (aravindg):

@Zekarias srry i was nt here